OpenStudy (anonymous):

find the sum of the infinite geometric series, if it exists. -30 -6 -6/5 -6/25 -.....

OpenStudy (anonymous):

is it times 5? do not exist is an answer choice

OpenStudy (anonymous):

does not exist*

OpenStudy (anonymous):

r is not 5, it's 1/5 (dividing by 5 = multiplying by 1/5). \[S_n=\frac{ a_1 }{ 1-r }\]

OpenStudy (anonymous):

a_1 is the first term and r is 1/5

OpenStudy (anonymous):

a_1=-30 a_2= -30/1-1/5?

OpenStudy (anonymous):

\(\frac{ -30 }{ 1-\frac{ 1 }{ 5 } }\) is the sum of the series. you don't need any term other than the first one. All you have to do is simplify that to one number

OpenStudy (anonymous):

37.5 = -75/2

OpenStudy (anonymous):

that's the answer

OpenStudy (anonymous):

ok cool thank you!

OpenStudy (anonymous):

you're welcome